Circular No. 7417

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000ce IN UGC 4195 T. Puckett, Mountain Town, GA, reports his discovery of an apparent

supernova (mag 15.0) on an unfiltered CCD frame (limiting mag 19.3) taken with the Puckett Observatory 0.35-m automated supernova patrol telescope on May 8.13 UT. SN 2000ce is located at α = 8^{h}05^{m}09^{s}.45, δ = +66^{o}47′15′′.2 (equinox 2000.0), which is 15′′.1 east and 17′′.3 north of the center of UGC 4195. The new object was also present on an unfiltered CCD frame taken on May 9.05, but it was not present on a frame taken on Mar. 23.23 or on Palomar Sky Survey images taken on 1997 Feb. 5 (limiting mag ∼ 21.0) or 1955 Jan. 29 (limiting mag ∼ 20.0).

VARIABLE STAR IN NGC 3432 R. M. Wagner, Ohio State University; G. D. Schmidt, P. Smith, and D.

Hines, Steward Observatory; and S. G. Starrfield, Arizona State University, report: “We obtained a spectrum (resolution 1.3 nm; range 432–740 nm) of the variable star in the field of NGC 3432 (cf. IAUC 7415) on May 6.2 UT with the Steward Observatory CCD SpectroPolarimeter attached to the Bok 2.3-m telescope at Kitt Peak. The spectrum exhibits relatively strong Balmer emission lines of Hα (equivalent width = 34.4 nm, FWHM = 1677 km/s) and Hβ (EW = 12.8 nm, FWHM = 2132 km/s) superposed on a smooth continuum. No other species are detected. The centroid of these emission lines yields a redshift of ∼ 0.002. This is consistent with the catalogued redshift of NGC 3432 (0.0020), indicating that the new variable is located in this galaxy and is not a foreground or background object. The spectral features resemble that of a classical nova observed in its early decline phase with an expansion velocity of ∼ 950 km/s. The V continuum magnitude inferred from our spectrum is 19.1 ± 0.3. If this variable reached an apparent magnitude at maximum of 17–18, then at the distance of NGC 3432, it would imply an absolute visual magnitude of at least −12, compared with the mean absolute magnitude of Galactic and M31 novae at maximum of ∼ −8. If this were a classical nova, then the peak bolometric magnitude, which implies a luminosity exceeding 10^{6} L_{⊙}, is inconsistent with the low expansion velocity that we measure.”

CI AQUILAE W. Liller, Viña del Mar, Chile, reports that nothing is present at the

position of CI Aql on 217 orange-red photographs (limiting mag typically ∼ 11.0) taken from Oct. 1983 to 2000 Apr. 4.333 UT.

2000 May 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green